Faced with two distinct demand functions:



Q1=24-0.2P1



Q2=10-0.05P2


1
Expert's answer
2022-01-19T14:24:04-0500

Faced with two distinct demand functions:


"Q_1=24-0.2P_1"

"Q_2=10-0.05P_2"

The average cost "AC=40+35\/Q." What price the firm will change with discrimination?


"Q_1=24-0.2P_1""P_1=\frac{24-Q_1}{0.2}"




"Q_2=10-0.05P_2""P_2=\frac{10-Q_2}{0.05}"


Average cost 

"AC=40+\frac{35}{Q}"


Total Revenue 

"TR=P\times Q"




"TR_1=P_1\times Q_1=\frac{24-Q_1}{0.2}\times Q_1"




"=120Q_1-5Q_1^2"




"TR_2=P_2\times Q_2=\frac{10-Q_2}{0.05}\times Q_2"




"=200Q_2-20Q_2^2"


Marginal revenue


"MR_1=\frac{d(TR_1)}{dQ_1}=120-10Q_1"




"MR_2=\frac{d(TR_2)}{dQ_2}=200-40Q_2"


Total cost

"TC=(40-\frac{35}{Q})\times Q=40Q-35"


Marginal cost


"MC=\frac{d(TC)}{dQ}=40"


Profit maximizing quantity of the firm will be when


"MR_1=MR_2=MC"




"MR_1=MC=>120-10Q_1=40"




"Q_1=8"

Then


"P_1=\frac{24-Q_1}{0.2}=\frac{24-8}{0.2}=80"




"MR_2=MC=>200-40Q_2=40"

"Q_2=4"


Then


"P_2=\frac{10-Q_2}{0.05}=\frac{10-4}{0.05}=120"



Therefore the prices firm will charge with discrimination are 

"P_1=80\space and\space P_2=120"